OpenStudy (anonymous):

HHHEEELLLPPP e^3x=10

OpenStudy (studygurl14):

are you solving for x? And I'm assuming the x is part of the exponent, right...?

OpenStudy (solomonzelman):

3x is the power, I would assume:) e^(3x)=10 hit both sides with a natural log, ln(e^(3x))=ln(10) 3xlne=ln10 3x=ln10 x=(ln10)/3

OpenStudy (studygurl14):

What do you know about euler's number?

OpenStudy (solomonzelman):

I actually alwyas forget that it is called that:)

OpenStudy (studygurl14):

Me too. I had to think for a little. :)